Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 7.4 pb-1.

Monday August 25
The day shift began with the Tevatron (TeV) in a squeeze, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 9:03 AM. At 10:51 AM, Operations established store 2943 with an initial luminosity of 4.0E31.

A Pbar expert began a Pbar study period at 11:14 AM.

At 11:45 AM, the I- Source was off due to a problem with magnet currents. Operators notified an expert who accessed the dome and discovered a bad power supply, which he replaced. The I- Source was back on line at 12:40 PM.

Pbar experts completed their studies and Pbar resumed stacking at 11:50 PM.

Tuesday August 26
The midnight shift began with Operation monitoring store 2943, with Pbar stacking, and with MiniBooNE taking beam.

At 3:45 AM, a Booster kicker power supply (B:MKS02) tripped off and wouldn't reset remotely. Operators responded, bypassed the alarm, and resumed stacking and beam to MiniBooNE at 4:03 AM.

At 4:27 AM, MKS02 tripped off due to sparking. Operators investigated and discovered a bad cable termination. Operations called in an expert and he had the kicker back up by 6 AM.

Operations terminated store 2943 at 6:16 AM.

A Main Injector (MI) expert began a two-hour study period at 6:20 AM.

A TeV expert began a study period at 8 AM.

A NuMI expert began at study period at 9:43 AM.

A Switchyard expert began a study period at 10 AM, but it didn't get very far. At 10:36 AM, a magnet (H203) showed a ground fault indication. An expert investigated and found that it was only a "sticky" switch. The study period began, again, at 11:26 AM.

At 12:23 PM, the TeV suffered a quench due to a lightening strike near the sector A3 service building. The strike actually caused a ground fault with the TeV electrical safety system with the quench resulting from that problem. There was no beam in the machine.

MiniBooNE asked to have their beam turned off for a short time at 1 PM due to data acquisition problems.

Operations established a TeV 36x0 study-store at 3:10 PM.

Operations established a TeV 36x4 study-store around 4:49 PM.

Operations put the TeV into shot setup at 6:30 PM.

Operations established store 2953 PM at 8:53 PM with an initial luminosity of 3.6E31. Pbar resumed stacking.

Earlier in the evening there were some complaints about domestic water flow. Domestic water is used mostly for sinks and toilets. The duty mechanic, thinking it was a bad pump, switched the domestic water supply to another well at 9:34 PM. However, domestic water flow remained off for the entire site. Water experts came in and at 11:35 PM and discovered at water leak between the Collider Detector Facility (CDF) and D-zero (D0). They isolated the leak from the rest of the site, but this left D0 and C0 still without water.

Wednesday August 27
The midnight shift began with MI experts conducting studies, with Pbar stacking, and with MiniBooNE taking beam.

A MI expert began a 2.5Mhz study at 5:41 AM. He was finished by 7:17 AM.

Today's Plan
The plan for today is to terminate store 2953 at 11 AM and then put the TeV into another shot setup.

Tonight, MI and Pbar will conduct 2.5Mhz transfers of antiprotons to the Recycler.

Tomorrow, August 28th, will be a shutdown day with the TeV off for about sixteen hours and the rest of the machine off for about eight hours.

The domestic water leak repairs will begin today and hopefully be done tonight or early tomorrow.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/27/2003   email Fermilab